The price elasticity for rice is estimated to be -0.4 and the income elasticity is 0.8. At a price of $0.40 per pound and a per capita income of $20,000, the demand for rice is 50 million tons per year.

a) if the price of rice increases to $0.41 per pound and income per capita remains at $20,000, what will be the quantity demanded?


1
Expert's answer
2022-03-17T14:03:22-0400

the change in quantity demanded is

"-0.4×\frac{(0.41 - 0.40)}{0.40}= -1"


so the new quantity demanded is

"Qd = 50 - 1 = 49" million tons